# Quantum Mycielskians: symmetries, twin vertices and distinguishing labelings

### v1

@@ 3. Quantum symmetries of $\mu_{r-1}(\mathcal{G})$ from quantum symmetries of $\mathcal{G}$
First, notice that $Q_{0}$ and $Q_{a}$ with $a=1,\ldots,n$ are projections. Indeed, we immediately notice that both $Q_{0}$ and $Q_{a}$ with $a=1,\ldots,n$ are idempotent. Furthermore, since $\iota_{k}$, $k=0,\ldots,r$ are $\ast$-maps, we see that all elements of $\mathfrak{P}_{\mu}^{r-1}$ are self-adjoint. Finally, we note that
@@
If $\rho_{\mathcal{G},\mathscr{H}}(P_{a})=P_{a}\otimes\mathds{1}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))}$, then the above expression simplifies into $Q_{a}\otimes\mathds{1}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))}$, so that $\Phi_{r,\mathscr{H}}$ induces a homomorphism
@@
such that $\Phi_{r,\mathscr{H}}^{\mathfrak{P}}=(\mathrm{id}\otimes\pi_{r}^{\mathfrak{P}})\circ\rho_{\scriptscriptstyle\mu_{r-1}(\mathcal{G}),\mathscr{H}}$. Moreover, $\pi_{r}^{\mathfrak{P}}$ restricts itself to a well-defined homomorphism $\pi_{r}^{\mathfrak{P},\mathfrak{P}_{\mu}^{r-1}}\colon\operatorname{C}(\operatorname{QAut}(\mu_{r-1}(\mathcal{G}))_{\mathfrak{P}_{\mu}^{r-1}})\to\operatorname{C}(\operatorname{QAut}(\mathcal{G})_{\mathfrak{P}})$ and this homomorphism is compatible with comultiplications (cf. [19, Proposition 4.7]) and is surjective by the last statement of Proposition 3.1. ∎
@@ 3.1. The role of twin vertices
Let $G=(V,E)$ be a classical graph. For $v\in V$ define $N(v)=\bigl{\{}u\in{V}\,\bigr{|}\bigl{.}\,\{u,v\}\in{E}\bigr{\}}$. Two vertices $x,y\in V$ are said to be twins if $N(x)=N(y)$.
Theorem 3.5 ([1, Proposition 4.1]).
@@
$A_{\mu(G)}=\begin{bmatrix}0&0&1\\ 0&A_{G}&A_{G}\\ 1&A_{G}&0\end{bmatrix}$
we deduce that $\rho_{12}=\rho_{21}=0$. Denoting $\rho_{i}=\rho_{ii}\in\operatorname{Aut}(G)$, for $i=1,2$, the remaining conditions are
$\rho_{2}A_{G}=A_{G}\rho_{1}=\rho_{1}A_{G}=A_{G}\rho_{2}.$ (3)
Now, evaluating both sides of the equation $\rho_{1}A_{G}=\rho_{2}A_{G}$ on $\delta_{v}$, the characteristic function of the vertex $v\in V$, we end up with equation $\rho_{1}\chi_{N(v)}=\rho_{2}\chi_{N(v)}$, where $\chi_{C}$ is the indicator function of the set $C$. Since $\rho_{1},\rho_{2}$ are automorphisms of the classical graph $G$, there exist permutations $\tau,\sigma$ of $V$ such that $\rho_{1}\chi_{N(v)}=\chi_{N(\tau(v))}$ and $\rho_{2}\chi_{N(v)}=\chi_{N(\sigma(v))}$. The lack of twin vertices implies that $\sigma=\tau$, so that $\rho_{1}=\rho_{2}$, i.e. every automorphism of $\mu(G)$ that has $\bullet$ as a fixed point is induced by an automorphism of $G$. ∎
@@
Note that in the proof presented above the conclusion $\rho_{1}=\rho_{2}$ can be immediately reached from (3) if we assume that $A_{G}$ is invertible. This is of course a stronger assumption than $G$ being twin-free.
@@
A graph $G=(V,E)$ has twin vertices if and only if its adjacency matrix $A_{G}$ has at least two same rows or, equivalently, columns. The existence of such two rows is equivalent to the existence of a permutation $\pi$ of the set $V$ such that $P_{\pi}A_{G}=A_{G}$, where $P_{\pi}$ is the corresponding permutation matrix. In contrast, the presence of the same columns is equivalent to the existence of $\sigma$ such that $A_{G}P_{\sigma}=A_{G}$.
@@
Let $\mathcal{G}=(\mathbb{V}_{\mathcal{G}},\psi_{\mathcal{G}},A_{\scriptscriptstyle\mathcal{G}})$ be a quantum graph and $\mathbb{H}$ be a compact quantum group. We say that $\mathcal{G}$ has $\mathbb{H}$-twin vertices, or simply $\mathbb{H}$-twins, if there exists a pair of different actions of $\mathbb{H}$ on $\mathcal{G}$
@@
Suppose that $\mathcal{G}$ and $\mathbb{H}$ are classical, i.e. $\mathbb{H}$ can be identified with a compact group $H$. Suppose that $\mathcal{G}$ has $H$-twins, i.e. there exist $\rho_{1},\rho_{2}\colon\operatorname{C}(\mathcal{G})\to\operatorname{C}(\mathcal{G})\otimes\operatorname{C}(\mathbb{H})$ satisfying (4). The actions $\rho_{1}$ and $\rho_{2}$ can be viewed as homomorphisms $H\ni h\mapsto\rho_{i}(h)\in\operatorname{Aut}(\mathcal{G})$. Since $\rho_{1}\neq\rho_{2}$ there is $h\in H$ such that $\rho_{1}(h)\neq\rho_{2}(h)$ and
@@
Defining $\sigma=\rho_{2}(h)^{-1}\rho_{1}(h)\in\operatorname{Aut}(\mathcal{G})$ we have $\sigma\circ A_{\scriptscriptstyle\mathcal{G}}=A_{\scriptscriptstyle\mathcal{G}}$. Since $\sigma\neq\mathrm{id}$ we can find two vertices $v_{1}\neq v_{2}$ and $\sigma(v_{1})=v_{2}$ and it is easy to see that $v_{1}$ and $v_{2}$ are twins.
Conversely, if $v_{1},v_{2}\in\mathbb{V}_{\mathcal{G}}$ are twins we can define an order two symmetry $\pi\in\operatorname{Aut}(\mathcal{G})$ such that $\pi(v)=v$ for all $v\in\mathbb{V}\setminus\{v_{1},v_{2}\}$ and $\pi(v_{1})=v_{2}$ (note that this is a different way of saying that $v_{1}$ and $v_{2}$ are twin vertices). Taking $\mathbb{H}=\mathbb{Z}_{2}=\{e,h\}$ and putting $\rho_{1}(h)=\pi$ and $\rho_{2}(h)=\mathrm{id}$ we see that $\mathcal{G}$ has $\mathbb{Z}_{2}$-twins in the sens of Definition 3.8.
Next, we generalize Theorem 3.5 to the case of quantum graphs.
@@
Let $\mathcal{G}=(\mathbb{V}_{\mathcal{G}},\psi_{\mathcal{G}},A_{\scriptscriptstyle\mathcal{G}})$ be a quantum graph that does not have quantum twin vertices. Let $\mathfrak{P}$ be a partition of unity in $\operatorname{C}(\mathcal{G})\otimes\operatorname{B}(\mathscr{H})$ with some (separable) Hilbert space $\mathscr{H}$. Then $\operatorname{QAut}(\mu(\mathcal{G}))_{\mathfrak{P}_{\mu}}\cong\operatorname{QAut}(\mathcal{G})_{\mathfrak{P}}$.
@@
By Corollary 3.3, $\operatorname{QAut}(\mathcal{G})_{\mathfrak{P}}$ is a closed quantum subgroup of $\operatorname{QAut}(\mu(\mathcal{G}))_{\mathfrak{P}_{\mu}}$. Denote the corresponding quotient map by $\pi_{2}^{\mathfrak{P},\mathfrak{P}_{\mu}}$. On the other hand, let
$\phi\colon\operatorname{C}(\mu(\mathcal{G}))\otimes\operatorname{B}(\mathscr{H})\longrightarrow\operatorname{C}(\mu(\mathcal{G}))\otimes\operatorname{B}(\mathscr{H})\otimes\operatorname{C}\bigl{(}\operatorname{QAut}(\mu(\mathcal{G}))_{\mathfrak{P}_{\mu}}\bigr{)}$
be the action of $\operatorname{QAut}(\mu(\mathcal{G}))_{\mathfrak{P}_{\mu}}$ on $\operatorname{C}(\mu(\mathcal{G}))\otimes\operatorname{B}(\mathscr{H})$.
Then $\phi=(\mathrm{id}\otimes\pi_{\mu})\rho_{\mu(\mathcal{G}),\mathscr{H}}$ with the quotient map $\pi_{\mu}\colon\operatorname{C}(\operatorname{QAut}(\mu(\mathcal{G})))\to\operatorname{C}(\operatorname{QAut}(\mu(\mathcal{G}))_{\mathfrak{P}_{\mu}})$. For $i,j\in\{0,1,2\}$, define $\phi_{i,j}\colon\operatorname{C}(\mathcal{G})\otimes\operatorname{B}(\mathscr{H})\to\operatorname{C}(\mathcal{G})\otimes\operatorname{B}(\mathscr{H})\otimes\operatorname{C}(\operatorname{QAut}(\mu(\mathcal{G}))_{\mathfrak{P}_{\mu}})$ by
@@
$\phi_{i,j}=(\iota_{i}\otimes\mathrm{id}_{\operatorname{B}(\mathscr{H})}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mu\circ(\mathcal{G}))_{\mathfrak{P}_{\mu}})})\circ\phi\circ\bigl{(}\iota_{j}^{\ast}\otimes\mathrm{id}_{\operatorname{B}(\mathscr{H})}\bigr{)}.$
It is easy to see that since $\left[\begin{smallmatrix}1\\ 0\\ 0\end{smallmatrix}\right]$ belongs to $\mathfrak{P}_{\mu}$ and $\left[\begin{smallmatrix}0\\ \mathds{1}\\ \mathds{1}\end{smallmatrix}\right]$ is a sum of elements of $\mathfrak{P}_{\mu}$, we have $\phi_{0,k}=0$ and $\phi_{k,0}=0$ for $k=1,2$, and $\phi_{0,0}(1)=1\otimes\mathds{1}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))}$. For $k,l=1,2$ let
$\widetilde{\phi}_{k,l}=(\mathrm{id}_{\operatorname{C}(\mathcal{G})}\otimes\pi_{2}^{\mathfrak{P},\mathfrak{P}_{\mu}})\phi_{k,l}\colon\operatorname{C}(\mathcal{G})\otimes\operatorname{B}(\mathscr{H})\to\operatorname{C}(\mathcal{G})\otimes\operatorname{B}(\mathscr{H})\otimes\operatorname{C}(\operatorname{QAut}(\mathcal{G})_{\mathfrak{P}}).$
Then
$\widetilde{\phi}_{k,l}=\bigl{(}\iota_{k}\otimes\mathrm{id}_{\operatorname{B}(\mathscr{H})}\otimes\Pi\bigr{)}\rho_{\mu(\mathcal{G})}\bigl{(}\iota_{l}^{\ast}\otimes\mathrm{id}_{\operatorname{B}(\mathscr{H})}\bigr{)},$
with $\Pi=\pi_{2}^{\mathfrak{P},\mathfrak{P}_{\mu}}\circ\pi_{\mu}$. Recall that
$A_{\mu(\mathcal{G})}=\delta_{\mathcal{G}}\iota_{2}\eta_{\mathcal{G}}\iota_{0}^{\ast}+\delta_{\mathcal{G}}\iota_{0}\eta_{\mathcal{G}}^{\ast}\iota_{2}^{\ast}+\iota_{1}A_{\mathcal{G}}\iota_{1}^{\ast}+\iota_{1}A_{\mathcal{G}}\iota_{2}^{\ast}+\iota_{2}A_{\mathcal{G}}\iota_{1}^{\ast},$
and $\rho_{\mu(\mathcal{G})}A_{\mu(\mathcal{G})}=\bigl{(}A_{\mu(\mathcal{G})}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mu(\mathcal{G})))}\bigr{)}\rho_{\mu(\mathcal{G})}$ which leads to
$\rho_{\mu(\mathcal{G}),\mathscr{H}}\bigl{(}A_{\mu(\mathcal{G})}\otimes\mathrm{id}_{\operatorname{B}(\mathscr{H})}\bigr{)}=\bigl{(}A_{\mu(\mathcal{G})}\otimes\mathrm{id}_{\operatorname{B}(\mathscr{H})}\otimes\mathrm{id}\bigr{)}\rho_{\mu(\mathcal{G}),\mathscr{H}}.$
Next, we explore the equation
$\bigl{(}\iota_{k}^{\ast}\otimes\mathrm{id}_{\operatorname{B}(\mathscr{H})}\otimes\Pi\bigr{)}$ $\rho_{\mu(\mathcal{G}),\mathscr{H}}(A_{\mu(\mathcal{G})}\otimes\mathrm{id}_{\operatorname{B}(\mathscr{H})})\bigl{(}\iota_{l}\otimes\mathrm{id}_{\operatorname{B}(\mathscr{H})}\bigr{)}$
@@
$=(\iota_{k}^{\ast}\otimes\mathrm{id}_{\operatorname{B}(\mathscr{H})}\otimes\Pi)\bigl{(}A_{\mu(\mathcal{G})}\otimes\mathrm{id}_{\operatorname{B}(\mathscr{H})}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mu(\mathcal{G})))}\bigr{)}\rho_{\mu(\mathcal{G}),\mathscr{H}}\bigl{(}\iota_{l}\otimes\mathrm{id}_{\operatorname{B}(\mathscr{H})}\bigr{)}$
for $k,l=0,1,2$. First, taking $(k,l)=(0,1)$ and $(k,l)=(1,0)$ leads to $\widetilde{\phi}_{2,1}=\widetilde{\phi}_{1,2}=0$. Using this and considering $(k,l)=(1,2)$ gives $\widetilde{\phi}_{1,1}A_{\mathcal{G}}=A_{\mathcal{G}}\widetilde{\phi}_{2,2}$, while $(k,l)=(2,1)$ leads to $\widetilde{\phi}_{2,2}A_{\mathcal{G}}=A_{\mathcal{G}}\widetilde{\phi}_{1,1}$. Finally, $(k,l)=(1,1)$ leads to $\widetilde{\phi}_{1,1}A_{\mathcal{G}}=A_{\mathcal{G}}\widetilde{\phi}_{1,1}$. Therefore, $\widetilde{\phi}_{1,1}A_{\mathcal{G}}=\widetilde{\phi}_{2,2}A_{\mathcal{G}}$. Since $\mathcal{G}$ does not have quantum twin vertices, we deduce $\widetilde{\phi}_{1,1}=\widetilde{\phi}_{2,2}$. This ends the proof. ∎

### v2

@@ 3. Quantum symmetries of $\mu_{r-1}(\mathcal{G})$ from quantum symmetries of $\mathcal{G}$
First, notice that $Q_{0}$ and $Q_{a}$ with $a=1,\ldots,n$ are projections. Indeed, we immediately notice that both $Q_{0}$ and $Q_{a}$ with $a=1,\ldots,n$ are idempotent and self-adjoint.Finally, we note that
@@
If $\rho_{\scriptscriptstyle\mathcal{G},\mathscr{H}}(P_{a})=P_{a}\otimes\mathds{1}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))}$, then the above expression simplifies into $Q_{a}\otimes\mathds{1}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))}$, so that $\Phi_{r,\mathscr{H}}$ induces a homomorphism
@@
such that $\Phi_{r,\mathscr{H}}^{\mathfrak{P}}=(\mathrm{id}\otimes\pi_{r}^{\mathfrak{P}})\circ\rho_{\scriptscriptstyle\mu_{r-1}(\mathcal{G}),\mathscr{H}}$. Moreover, $\pi_{r}^{\mathfrak{P}}$ restricts itself to a well-defined homomorphism $\pi_{r}^{\mathfrak{P},\mathfrak{P}_{\mu}^{r-1}}\colon\operatorname{C}(\operatorname{QAut}(\mu_{r-1}(\mathcal{G}))_{\mathfrak{P}_{\mu}^{r-1}})\to\operatorname{C}(\operatorname{QAut}(\mathcal{G})_{\mathfrak{P}})$ and this homomorphism is compatible with comultiplications (cf. [22, Proposition 4.7]) and is surjective by the last statement of Proposition 3.1. ∎
@@ 3.1. The role of twin vertices
Let $G=(V,E)$ be a classical graph. For $v\in V$ define $N(v)=\bigl\{u\in{V}\,\bigr|\bigl.\,\{u,v\}\in{E}\bigr\}$, the neighbourhood of $v$. Two vertices $x,y\in V$ are said to be twins if $N(x)=N(y)$.
Theorem 3.5 ([2, Proposition 4.1]).
@@
$A_{\scriptscriptstyle\mu(G)}=\begin{bmatrix}0&0&1\\ 0&A_{\scriptscriptstyle G}&A_{\scriptscriptstyle G}\\ 1&A_{\scriptscriptstyle G}&0\end{bmatrix}$
[…]
$\rho_{2}A_{\scriptscriptstyle G}=A_{\scriptscriptstyle G}\rho_{1}=\rho_{1}A_{\scriptscriptstyle G}=A_{\scriptscriptstyle G}\rho_{2}.$ (3)
Now, evaluating both sides of the equation $\rho_{1}A_{\scriptscriptstyle G}=\rho_{2}A_{\scriptscriptstyle G}$ on $\delta_{v}$, the characteristic function of the vertex $v\in V$, we end up with equation $\rho_{1}\chi_{N(v)}=\rho_{2}\chi_{N(v)}$, where $\chi_{C}$ is the indicator function of the set $C$. Since $\rho_{1},\rho_{2}$ are automorphisms of the classical graph $G$, there exist permutations $\tau,\sigma$ of $V$ such that $\rho_{1}\chi_{N(v)}=\chi_{N(\tau(v))}$ and $\rho_{2}\chi_{N(v)}=\chi_{N(\sigma(v))}$. The lack of twin vertices implies that $\sigma=\tau$, so that $\rho_{1}=\rho_{2}$, i.e. every automorphism of $\mu(G)$ that has $\bullet$ as a fixed point is induced by an automorphism of $G$. ∎
@@
Note that in the proof presented above the conclusion $\rho_{1}=\rho_{2}$ can be immediately reached from (3) if we assume that $A_{\scriptscriptstyle G}$ is invertible. This is of course a stronger assumption than $G$ being twin-free.
@@
A graph $G=(V,E)$ has twin vertices if and only if its adjacency matrix $A_{\scriptscriptstyle G}$ has at least two same rows or, equivalently, columns. The existence of such two rows is equivalent to the existence of a permutation $\pi$ of the set $V$ such that $P_{\pi}A_{\scriptscriptstyle G}=A_{\scriptscriptstyle G}$, where $P_{\pi}$ is the corresponding permutation matrix. In contrast, the presence of the same columns is equivalent to the existence of $\sigma$ such that $A_{\scriptscriptstyle G}P_{\sigma}=A_{\scriptscriptstyle G}$.
@@
Let $\mathcal{G}=(\mathbb{V}_{\scriptscriptstyle\mathcal{G}},\psi_{\scriptscriptstyle\mathcal{G}},A_{\scriptscriptstyle\mathcal{G}})$ be a quantum graph and $\mathbb{H}$ be a compact quantum group. We say that $\mathcal{G}$ has $\mathbb{H}$-twin vertices, or simply $\mathbb{H}$-twins, if there exists a pair of different actions of $\mathbb{H}$ on $\mathcal{G}$
@@
Suppose that $\mathcal{G}$ and $\mathbb{H}$ are classical, i.e. $\mathbb{H}$ can be identified with a compact group $H$. Suppose that $\mathcal{G}$ has $H$-twins, i.e. there exist $\rho_{1},\rho_{2}\colon\operatorname{C}(\mathcal{G})\to\operatorname{C}(\mathcal{G})\otimes\operatorname{C}(H)$ satisfying (4). The actions $\rho_{1}$ and $\rho_{2}$ can be viewed as homomorphisms $H\ni h\mapsto\rho_{i}(h)\in\operatorname{Aut}(\mathcal{G})$. Since $\rho_{1}\neq\rho_{2}$ there is $h\in H$ such that $\rho_{1}(h)\neq\rho_{2}(h)$ and
@@
Defining $\sigma=\rho_{2}(h)^{-1}\rho_{1}(h)\in\operatorname{Aut}(\mathcal{G})$ we have $P_{\sigma}A_{\scriptscriptstyle\mathcal{G}}=A_{\scriptscriptstyle\mathcal{G}}$. Since $\sigma\neq\mathrm{id}$ we can find two vertices $v_{1}\neq v_{2}$ and $\sigma(v_{1})=v_{2}$ and it is easy to see that $v_{1}$ and $v_{2}$ are twins.
Conversely, if $v_{1},v_{2}\in\mathbb{V}_{\mathcal{G}}$ are twins we can define an order two symmetry $\pi\in\operatorname{Aut}(\mathcal{G})$ such that $\pi(v)=v$ for all $v\in\mathbb{V}\setminus\{v_{1},v_{2}\}$ and $\pi(v_{1})=v_{2}$ (note that this is a different way of saying that $v_{1}$ and $v_{2}$ are twin vertices). Taking $\mathbb{H}=\mathbb{Z}_{2}=\{e,h\}$ and putting $\rho_{1}(h)=\pi$ and $\rho_{2}(h)=\mathrm{id}$ we see that $\mathcal{G}$ has $\mathbb{Z}_{2}$-twins in the sense of Definition 3.8.
Next, we generalize Theorem 3.5 to the case of quantum graphs, with a possible amplification by $\operatorname{B}(\mathscr{H})$. In the next Theorem we use partitions of unity and the corresponding construction of $\mathfrak{P}_{\mu}$ as described in Lemma 3.2 and the notation below. In what follows $\operatorname{K}(\mathscr{H})$ denoted the algebra of compact operators on the Hilbert space $\mathscr{H}$ and $\operatorname{M}(\cdot)$ is the multiplier functor. The reader will find more on these concepts and their use in introductory sections of [25] and [26].
@@
Let $\mathcal{G}=(\mathbb{V}_{\scriptscriptstyle\mathcal{G}},\psi_{\scriptscriptstyle\mathcal{G}},A_{\scriptscriptstyle\mathcal{G}})$ be a quantum graph that does not have quantum twin vertices. Let $\mathfrak{P}$ be a partition of unity in $\operatorname{C}(\mathcal{G})\otimes\operatorname{B}(\mathscr{H})$ with some (separable) Hilbert space $\mathscr{H}$. Then $\operatorname{QAut}(\mu(\mathcal{G}))_{\mathfrak{P}_{\mu}}\cong\operatorname{QAut}(\mathcal{G})_{\mathfrak{P}}$.
@@
Let us first recall that $\operatorname{QAut}(\mathcal{G})_{\mathfrak{P}}$ can be viewed as a closed quantum subgroup of $\operatorname{QAut}(\mu(\mathcal{G}))_{\mathfrak{P}_{\mu}}$ (cf. Corollary  3.3). We shall prove that the two quantum groups are in fact identical.
So let us consider a partition of unity $\mathfrak{P}\subset\operatorname{C}(\mathcal{G})\otimes\operatorname{B}(\mathscr{H})$. Note that the partition $\mathfrak{P}_{\mu}\subset\operatorname{C}(\mu(\mathcal{G}))\otimes\operatorname{B}(\mathscr{H})$ assigned to $\mathfrak{P}$ contains the projection $P_{\bullet}\otimes\mathds{1}_{\operatorname{B}(\mathscr{H})}\in\operatorname{C}(\mu(\mathcal{G}))\otimes\operatorname{B}(\mathscr{H})$ where $P_{\bullet}=\left[\begin{smallmatrix}1\\ 0\\ 0\end{smallmatrix}\right]\in\operatorname{C}(\mu(\mathcal{G}))$. Let $\rho_{\scriptscriptstyle\mu(\mathcal{G}),\mathscr{H}}$ be the amplification of the canonical action of $\operatorname{QAut}(\mu(\mathcal{G}))_{\mathfrak{P}_{\mu}}$ on $\operatorname{C}(\mu(\mathcal{G}))$. The direct sum structure of $\operatorname{C}(\mu(\mathcal{G}))$ gives rise to the “matrix structure” of $\rho_{{\scriptscriptstyle\mu(\mathcal{G}),\mathscr{H}}}=[\rho_{ij}]_{i,j=0,1,2}$ where using notation $\operatorname{C}(\mathbb{X}_{0})=\mathbb{C}$, $\operatorname{C}(\mathbb{X}_{1})=\operatorname{C}(\mathcal{G})$ and $\operatorname{C}(\mathbb{X}_{2})=\operatorname{C}(\mathcal{G})$ we define $\rho_{ij}:\operatorname{C}(\mathbb{X}_{j})\otimes\operatorname{K}(\mathscr{H})\to\operatorname{M}(\operatorname{C}(\mathbb{X}_{i})\otimes\operatorname{K}(\mathscr{H})\otimes\operatorname{C}(\operatorname{QAut}(\mu(\mathcal{G}))_{\mathfrak{P}_{\mu}})$ by the formula
@@
$\rho_{ij}=\bigl(\iota_{j}^{*}\otimes\mathrm{id}_{\operatorname{B}(\mathscr{H})}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mu(\mathcal{G}))_{\mathfrak{P}_{\mu}})}\bigr)\circ\rho_{{\scriptscriptstyle\mu(\mathcal{G}),\mathscr{H}}}\circ\bigl(\iota_{i}\otimes\mathrm{id}_{\operatorname{B}(\mathscr{H})}\bigr).$
Remembering that
@@
$\rho_{{\scriptscriptstyle\mu(\mathcal{G}),\mathscr{H}}}\circ\iota_{0}(\lambda)=\sqrt{1+2\delta_{\scriptscriptstyle\mathcal{G}}^{2}}\,\lambda\bigl(P_{\bullet}\otimes\mathds{1}_{\operatorname{B}(\mathscr{H})}\otimes\mathds{1}_{\operatorname{C}(\operatorname{QAut}(\mu(\mathcal{G}))_{\mathfrak{P}_{\mu}})}\bigr)$
for every $\lambda\in\mathbb{C}$ and using $\iota_{j}^{*}(P_{\bullet})=0$ for $j\in\{1,2\}$ we get $\rho_{j0}=0$ and and $\rho_{00}(\lambda)=\lambda\bigl(P_{\bullet}\otimes\mathds{1}_{\operatorname{B}(\mathscr{H})}\otimes\mathds{1}_{\operatorname{C}(\operatorname{QAut}(\mu(\mathcal{G}))_{\mathfrak{P}_{\mu}})}\bigr)$ for every $\lambda\in\mathbb{C}$. Using the *-condition of the *-homomorphism $\rho_{\scriptscriptstyle\mu(\mathcal{G}),\mathscr{H}}$ we get $\rho_{01}=0=\rho_{02}$.
Having the above facts established we can now prove that $\rho_{12}=\rho_{21}=0$. Indeed, since the adjacency matrix $A_{\scriptscriptstyle\mu(\mathcal{G})}$ is of the form
$A_{\scriptscriptstyle\mu(\mathcal{G})}=\delta_{\scriptscriptstyle\mathcal{G}}\iota_{2}\eta_{\scriptscriptstyle\mathcal{G}}\iota_{0}^{\ast}+\delta_{\scriptscriptstyle\mathcal{G}}\iota_{0}\eta_{\scriptscriptstyle\mathcal{G}}^{\ast}\iota_{2}^{\ast}+\iota_{1}A_{\scriptscriptstyle\mathcal{G}}\iota_{1}^{\ast}+\iota_{1}A_{\scriptscriptstyle\mathcal{G}}\iota_{2}^{\ast}+\iota_{2}A_{\scriptscriptstyle\mathcal{G}}\iota_{1}^{\ast}.$
(cf. Section 1.4) we get
$A_{\scriptscriptstyle\mu(\mathcal{G})}\bigl(\iota_{0}(1)\bigr)=\delta_{\scriptscriptstyle\mathcal{G}}\iota_{2}(\mathds{1}_{\operatorname{C}(\mathcal{G})})$
and therefore
$\bigl(\iota_{1}^{*}\otimes\mathrm{id}_{\operatorname{B}(\mathscr{H})}\otimes\mathrm{id}_{\operatorname{C}(\mu(\mathcal{G})}\bigr)\rho_{{\scriptscriptstyle\mu(\mathcal{G}),\mathscr{H}}}\bigl(A_{\scriptscriptstyle\mu(\mathcal{G})}(\iota_{0}(1))\otimes\mathds{1}_{\operatorname{B}(\mathscr{H})}\bigr)=\delta_{\scriptscriptstyle\mathcal{G}}\rho_{12}\bigl(\mathds{1}_{\operatorname{C}(\mathcal{G})}\otimes\mathds{1}_{\operatorname{B}(\mathscr{H})}\bigr).$
A similar computation but with the amplified adjacency matrix $A_{\scriptscriptstyle\mu(\mathcal{G})}$ to the left of the action $\rho_{\scriptscriptstyle\mu(\mathcal{G}),\mathscr{H}}$ shows that the left hand side of the above equality vanishes, and consequently $\rho_{12}(\mathds{1})=0$. This implies that $\rho_{12}=\rho_{21}=0$.
Using the vanishing of $\rho_{ij},\rho_{0j},\rho_{i0}$ for $i,j\in\{1,2\}$ proven so far and the equality
$\rho_{{\scriptscriptstyle\mu(\mathcal{G}),\mathscr{H}}}\bigl(A_{\scriptscriptstyle\mu(\mathcal{G})}\iota_{2}(x)\otimes\mathds{1}_{\operatorname{B}(\mathscr{H})}\bigr)=\bigl(A_{\scriptscriptstyle\mu(\mathcal{G})}\otimes\mathrm{id}_{\operatorname{K}(\mathscr{H})}\otimes\mathrm{id}_{\operatorname{C}(\mu(\mathcal{G}))}\bigr)\rho_{{\scriptscriptstyle\mu(\mathcal{G}),\mathscr{H}}}\bigl(\iota_{2}(x)\otimes\mathds{1}_{\operatorname{B}(\mathscr{H})}\bigr)$
which holds for every $x\in\operatorname{C}(\mathcal{G})$ we easily conclude that $\rho_{11}(A_{\scriptscriptstyle\mathcal{G}}\otimes\mathds{1}_{\operatorname{B}(\mathscr{H})})=\rho_{22}(A_{\scriptscriptstyle\mathcal{G}}\otimes\mathds{1}_{\operatorname{B}(\mathscr{H})})$. “Deamplifying” the later equality an referring to the assumed lack of quantum twins in $\mathcal{G}$ we conclude that $\rho_{11}=\rho_{22}$.
Summarizing we have so far proved that the canonical action of $\operatorname{QAut}(\mu(\mathcal{G}))_{\mathfrak{P}_{\mu}}$ on $\operatorname{C}(\mu(\mathcal{G}))$ is induced (in the sense described in Section 3) by the action of $\operatorname{QAut}(\mathcal{G})$. Moreover using the fact that $\rho_{\scriptscriptstyle\mu(\mathcal{G})}$ preserves $\mathfrak{P}_{\mu}$ we see that in fact it is induced by the action of $\operatorname{QAut}(\mathcal{G})_{\mathfrak{P}}$ and therefore we conclude that $\operatorname{QAut}(\mathcal{G})_{\mathfrak{P}}$ and $\operatorname{QAut}(\mu\mathcal{G})_{\mathfrak{P}_{\mu}}$ are indeed isomorphic. ∎